# Large Language Model Aided QoS Prediction for Service Recommendation

### v2

@@ I Introduction
The rapid development of cloud computing has led to an explosive growth in the number of cloud services, resulting in many services having identical or similar functionalities [1]. There is a growing need for service providers to provide the most suitable service for users to meet customer demands. When there exist many candidate services for a user request, quality of service (QoS) is a commonly used metric to measure the performance of these services. QoS refers to non-functionality aspects of services [2] such as throughput, response time, and cost. It plays a crucial role in distinguishing similar services and helps users select the most appropriate one. However, in real-world scenarios, evaluating the QoS of a large number of services on the client side is expensive and time-consuming due to the sheer scale. Additionally, most end users lack the expertise to do the evaluation [3]. Therefore, it is common for an end user is to predict the QoS values for an unknown service, and the QoS values are used to help the user find the most suitable service.
The problem of QoS prediction is illustrated in Figure I. For the sets of web users (u1 $\sim$ u4) and services (s1 $\sim$ s4) exists a QoS matrix, in which each value corresponds to a user-service interaction. The matrix contains some known QoS values from historical user-service interactions, and our goal is to predict unknown QoS values (marked by “?”) using these known values, thereby recommending more suitable services to users. Many methods have been developed to predict the QoS values for a given user and service. Among them, collaborative filtering is the most widely used technique. It aim to accurately estimate unknown QoS values by learning from the QoS value matrix from historical user-service interactions. However, in practice, the historical QoS value matrix is often highly sparse, as each user usually only interacted with a limited number of cloud services in the past. This data sparsity issue poses difficulty for many QoS prediction methods. Although many methods aim to mitigate this difficulty, it is still challenging to achieve ideal predictive performance.
@@ II-A QoS Prediction for Service Recommendation
Deep learning technology has demonstrated remarkable capabilities across various fields of artificial intelligence in recent years [23, 24, 25, 4, 5]. And an increasing number of deep learning based techniques for QoS prediction are developed [26, 27] too. Zou et al. [28] propose a novel adaptive QoS prediction framework, utilizing a dual-tower deep residual network to extract latent features of users and services and predict QoS values. Wu et al. [29] introduce a reputation-integrated graph convolutional network for robust and accurate QoS prediction. Wei et al. [30] use a transformer based method to learn the graph structure for time-aware service recommendation. Zhu et al. [31] develop a twin graph network based on graph contrastive learning, effectively addressing cold-start and data sparsity issues. Zhou et al. [32] propose the spatial context-aware time series forecasting (SCATSF) framework by utilizing both temporal and spatial contextual information of users and services. Lian et al. [33] facilitate multi-task learning for training QoS prediction model.
@@ IV Experiments
Gain 11.36% 3.84% 16.72% 1.41% 18.93% 2.01% 21.40% 0.21%
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
@@ IV-A Dataset
To evaluate the performance of our model, we conducted extensive experiments on the WSDream dataset [49]. WSDream is a large-scale real-world dataset of Internet user-service interactions that involves 5,825 services and 339 users in total. Each user and service is identified by a unique ID number. In addition to the IDs, WSDream also provides basic attribute information for each user and service. User attributes consist of IP Address, Country, IP Number, Autonomous System, Latitude, and Longitude. Service attributes include WSDL Address, Service Provider, IP Address, Country, IP Number, Autonomous System, Latitude, and Longitude. We utilize both the IDs and attributes for the users and services in our llmQoS model.
WSDream contains QoS record that include 1,873,850 response time values and 1,831,265 throughput values. For experiments, both the throughput sub-dataset and the response time sub-dataset are randomly split into training and testing sets. We use different density values $D$, which is the percentage of training set data in the total dataset, to test the models’ prediction ability at different data sparsity levels. Specifically we use $D$=5%, 10%, 15%, and 20%. It is expected that with all other settings unchanged, QoS prediction models achieve higher precision with larger values of $D$.
@@ IV-C Baseline Models
In our experiment, we selected 7 baselines to compare with our model and evaluate its performance. This includes traditional algorithms and the latest research. We keep dataset split method and evaluation settings for all baselines same to the proposed llmQoS for fair comparison.
UIPCC [18] is a traditional domain-based collaborative filtering method that predicts missing QoS values by leveraging historical invocation information of users and services.
@@
LACF [57] is a location-aware collaborative filtering method that integrates user and service location information to recommend web services to users.
PMF [58] is an enhanced traditional matrix factorization method that incorporates probabilistic factors into matrix decomposition for QoS prediction.
@@
LMF-PP [59] is a location-based matrix factorization method that integrates invocation and domain information, and utilizes fused similarity for preference propagation, addressing cold-start issues to achieve QoS prediction.
DCALF [60] detects domains and noise of users and services from the original QoS matrix, introducing density peak-based clustering methods during modeling for simultaneous detection of neighborhood and noise, achieving QoS prediction.
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
@@
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
@@
(a) Throughput $D$=5%
(b) Throughput $D$=10%
(c) Throughput $D$=15%
(d) Throughput $D$=20%
(e) Response time $D$=5%
(f) Response time $D$=10%
(g) Response time $D$=15%
(h) Response time $D$=20%
@@
FIGURE VIII: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model different depth and width configurations for the MLP network (§ III-B). Depth refers to the number of layers in the MLP. The main experiments use 3 layers. For different width configurations, we multiply the dimensionalities of each layers used by the main experiments by different width factor for comparison. We calculate and show the relative change of MAE and RMSE of a depth or width configuration from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. Generally, the effect of configuration of MLP has on the QoS prediction performance is insignificant, with changes less than 5% in most cases.
@@
FIGURE IX: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model trained using different batch sizes (BS) and learning rates (LR). We calculate and show the relative change of MAE and RMSE of a batch size or learning rate from the main experiments. This experiment is conducted with density level $D$=5and using Phi3mini. The performance is generally insensitive to batch size. It is sensitive to learning rate, with the optimal value around $10^{-4}$.
@@ IV-D Comparison with Baselines
The performance of the proposed llmQoS is compared with the baseline methods in Table I. Phi3mini is used for llmQoS in this experiment. We compare both MAE and RMSE for throughput and response time at different densities. It is evident that llmQoS outperforms all baselines consistently at all densities. For throughput value prediction, llmQoS can reduce the MAE by more than 20%, and RMSE by more than 6% from the best baseline model. For response time value prediction, llmQoS can also reduce MAE for more than 10% from the best baseline method. The results demonstrate that with the rich features from the descriptive sentences extracted by LLMs, the QoS prediction model can effectively find the similarity and correlation among users and services. Along with the information learned from historical interactions, llmQoS can achieve high QoS prediction accuracy.
@@ IV-G Ablation Study Using Random Feature
To verify this, we train a model that still has 3,072-dimensional feature vectors as input. But those feature vectors are randomly initialized for each user and service, so they do not contain any useful information. However, this model has the same number of parameters as the llmQoS model with Phi3mini feature. We compare the QoS prediction performance of this model with both the ID only model as in § IV-F and the llmQoS with Phi3mini in Figure VI. Generally, the model with random feature does not outperform ID only model. On throughput dataset under $D$=5and $D$=15%, the accuracy is marginally higher than the ID only model. But under other settings the performance actually worsens. We also observe that with random feature, the model converges with much less epochs than using Phi3mini feature. This indicates that more trainable parameters lead to overfitting. We verify that the performance boost of llmQoS is from the useful information encoded in the LLM feature, instead of added parameters of the model.
@@ IV-I Hyper-Parameter Tuning of MLP
We test the performance of llmQoS using different configurations for the MLP model as described in § III-B. For this experiment we test the density level $D$=5%, and using Phi3mini as the LLM feature extractor. We vary the depth of the MLP by adding or removing layers from it. The QoS prediction performance of different configurations is shown in 8(a) and 8(b). We change the width of the MLP by multiplying a width factor to the dimensionalities of the layers in it. The QoS prediction performance of different configurations is shown in 8(c) and 8(d).
@@ IV-J Training Parameter Tuning
We test the performance of llmQoS trained using different batch sizes and learning rates. For this experiment we test the density level $D$=5%, and using Phi3mini as the LLM feature extractor. The same configuration of MLP used in the main experiments is used.

### v3

@@ I Introduction
The rapid development of cloud computing has led to an explosive growth in the number of cloud services, resulting in many services having identical or similar functionalities [1]. There is a growing need for service providers to provide the most suitable service for users to meet customer demands. When there exist many candidate services for a user request, quality of service (QoS) is a commonly used metric to measure the performance of these services. QoS refers to non-functionality aspects of services [2, 3, 4] such as throughput, response time, and cost. It plays a crucial role in distinguishing similar services and helps users select the most appropriate one. However, in real-world scenarios, evaluating the QoS of a large number of services on the client side is expensive and time-consuming due to the sheer scale. Additionally, most end users lack the expertise to do the evaluation [5]. Therefore, it is common for an end user is to predict the QoS values for an unknown service, and the QoS values are used to help the user find the most suitable service.
The problem of QoS prediction is illustrated in Figure I. For the sets of web users (u1 $\sim$ u4) and services (s1 $\sim$ s4) exists a QoS matrix, in which each value corresponds to a user-service interaction. The matrix contains some known QoS values from historical user-service interactions, and our goal is to predict unknown QoS values (marked by “?”) using these known values, thereby recommending more suitable services to users. Many methods have been developed to predict the QoS values for a given user and service. Among them, collaborative filtering is the most widely used technique. It aim to accurately estimate unknown QoS values by learning from the QoS value matrix from historical user-service interactions. However, in practice, the historical QoS value matrix is often highly sparse [6] , as each user usually only interacted with a limited number of cloud services in the past. This data sparsity issue poses difficulty for many QoS prediction methods. Although many methods aim to mitigate this difficulty, it is still challenging to achieve ideal predictive performance.
@@ II-A QoS Prediction for Service Recommendation
Deep learning technology has demonstrated remarkable capabilities across various fields of artificial intelligence in recent years [26, 27, 28, 7, 29, 8]. And an increasing number of deep learning based techniques for QoS prediction are developed [30, 31, 32] too. Zou et al. [33] propose a novel adaptive QoS prediction framework, utilizing a dual-tower deep residual network to extract latent features of users and services and predict QoS values. Wu et al. [34] introduce a reputation-integrated graph convolutional network for robust and accurate QoS prediction. Wei et al. [35] use a transformer based method to learn the graph structure for time-aware service recommendation. Zhu et al. [36] develop a twin graph network based on graph contrastive learning, effectively addressing cold-start and data sparsity issues. Zhou et al. [37] propose the spatial context-aware time series forecasting (SCATSF) framework by utilizing both temporal and spatial contextual information of users and services. Lian et al. [38] facilitate multi-task learning for training QoS prediction model.
@@ IV Experiments
Gain 11.36% 3.84% 16.72% 1.41% 18.93% 2.01% 21.40% 0.19%
(a) Throughput DD=5%
(b) Throughput DD=10%
(c) Throughput DD=15%
(d) Throughput DD=20%
(e) Response time DD=5%
(f) Response time DD=10%
(g) Response time DD=15%
(h) Response time DD=20%
@@ IV-A Dataset
To evaluate the performance of our model, we conducted extensive experiments on the WSDream dataset [54]. WSDream is a large-scale real-world dataset of Internet user-service interactions, and is widely recognized as a standard benchmark for evaluating QoS prediction models in this research domain, involving 5,825 services and 339 users in total. Each user and service is identified by a unique ID number. In addition to the IDs, WSDream also provides basic attribute information for each user and service. User attributes consist of IP Address, Country, IP Number, Autonomous System, Latitude, and Longitude. Service attributes include WSDL Address, Service Provider, IP Address, Country, IP Number, Autonomous System, Latitude, and Longitude. We utilize both the IDs and attributes for the users and services in our llmQoS model.
WSDream contains QoS record that include 1,873,850 response time values and 1,831,265 throughput values. For experiments, both the throughput sub-dataset and the response time sub-dataset are randomly split into training and testing sets. We use different density values $D$, which is the percentage of training set data in the total dataset, to test the models’ prediction ability at different data sparsity levels. Specifically we use DD=5%, 10%, 15%, and 20%. It is expected that with all other settings unchanged, QoS prediction models achieve higher precision with larger values of $D$.
@@ IV-C Baseline Models
We benchmark llmQoS against several diverse QoS prediction baselines, which span a range of methodologies from traditional approaches like classic collaborative filtering and matrix factorization, to location-aware techniques, and more recent graph-based or optimization methods. All baselines use identical experimental settings for fair comparison.
UIPCC [21] is a classic collaborative filtering method predicting QoS from historical user-service invocation data.
@@
LACF [62] is a location-aware collaborative filtering method using integrated user and service location data for QoS prediction.
PMF [63] is a probabilistic matrix factorization technique applied to QoS prediction.
@@
PSO-USRec [64] is an optimization-based QoS prediction model using particle swarm optimization to search the global QoS distribution space and refine predictions.
LMF-PP [65] is a location-based matrix factorization method integrating invocation/domain information and fused similarity for preference propagation in QoS prediction.
DCALF [66] is a method to detect domains and noise in the QoS matrix via density peak-based clustering for QoS prediction.
(a) Throughput DD=5%
(b) Throughput DD=10%
(c) Throughput DD=15%
(d) Throughput DD=20%
(e) Response time DD=5%
(f) Response time DD=10%
(g) Response time DD=15%
(h) Response time DD=20%
@@
(a) Throughput DD=5%
(b) Throughput DD=10%
(c) Throughput DD=15%
(d) Throughput DD=20%
(e) Response time DD=5%
(f) Response time DD=10%
(g) Response time DD=15%
(h) Response time DD=20%
@@
(a) Throughput DD=5%
(b) Throughput DD=10%
(c) Throughput DD=15%
(d) Throughput DD=20%
(e) Response time DD=5%
(f) Response time DD=10%
(g) Response time DD=15%
(h) Response time DD=20%
@@
FIGURE VIII: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model different depth and width configurations for the MLP network (§ III-B). Depth refers to the number of layers in the MLP. The main experiments use 3 layers. For different width configurations, we multiply the dimensionalities of each layers used by the main experiments by different width factor for comparison. We calculate and show the relative change of MAE and RMSE of a depth or width configuration from the main experiments. This experiment is conducted with density level DD=5% and using Phi3mini. Generally, the effect of configuration of MLP has on the QoS prediction performance is insignificant, with changes less than 5% in most cases.
@@
FIGURE IX: QoS prediction performance in terms Mean Average Error (MAE) and Root Mean Squared Error (RMSE) of the proposed llmQoS model trained using different batch sizes (BS) and learning rates (LR). We calculate and show the relative change of MAE and RMSE of a batch size or learning rate from the main experiments. This experiment is conducted with density level DD=5% and using Phi3mini. The performance is generally insensitive to batch size. It is sensitive to learning rate, with the optimal value around $10^{-4}$.
@@ IV-D Comparison with Baselines
The performance of the proposed llmQoS model is compared with the baseline methods in Table I. Phi3mini is used for llmQoS in this experiment. We compare both MAE and RMSE for throughput and response time at different densities. It is evident that llmQoS outperforms all baselines consistently at all densities. For throughput value prediction, llmQoS can reduce the MAE by more than 20%, and RMSE by more than 6% from the best baseline model. For response time value prediction, llmQoS can also reduce MAE for more than 10% from the best baseline method. The prediction accuracy of every method all increases as the QoS matrix density grows. This indicates that with more observed data QoS values, CF-based methods can calculate similarities more accurately, and model-based approaches can capture additional latent features among users and services, leading to better QoS prediction. The results demonstrate that with the rich features from the descriptive sentences extracted by LLMs, the QoS prediction model can effectively find the similarity and correlation among users and services. Along with the information learned from historical interactions, llmQoS can achieve high QoS prediction accuracy.
@@ IV-G Ablation Study Using Random Feature
To verify this, we train a model that still has 3,072-dimensional feature vectors as input. But those feature vectors are randomly initialized for each user and service, so they do not contain any useful information. However, this model has the same number of parameters as the llmQoS model with Phi3mini feature. We compare the QoS prediction performance of this model with both the ID only model as in § IV-F and the llmQoS with Phi3mini in Figure VI. Generally, the model with random feature does not outperform ID only model. On throughput dataset under DD=5% and DD=15%, the accuracy is marginally higher than the ID only model. But under other settings the performance actually worsens. We also observe that with random feature, the model converges with much less epochs than using Phi3mini feature. This indicates that more trainable parameters lead to overfitting. We verify that the performance boost of llmQoS is from the useful information encoded in the LLM feature, instead of added parameters of the model.
@@ IV-I Hyper-Parameter Tuning of MLP
We test the performance of llmQoS using different configurations for the MLP model as described in § III-B. For this experiment we test the density level DD=5%, and using Phi3mini as the LLM feature extractor. We vary the depth of the MLP by adding or removing layers from it. The QoS prediction performance of different configurations is shown in 8(a) and 8(b). We change the width of the MLP by multiplying a width factor to the dimensionalities of the layers in it. The QoS prediction performance of different configurations is shown in 8(c) and 8(d).
@@ IV-J Training Parameter Tuning
We test the performance of llmQoS trained using different batch sizes and learning rates. For this experiment we test the density level DD=5%, and using Phi3mini as the LLM feature extractor. The same configuration of MLP used in the main experiments is used.
@@
IV-K Limitations
In various experiments, we have shown that the proposed llmQoS model has promising QoS prediction capabilities. However, our study has limitations. The straightforward MLP architecture of the llmQoS model, while demonstrating effectiveness, may not capture all deeply complex user-service interaction patterns. The computational overhead associated with LLMs remains also a practical consideration. The impact on QoS prediction performance of the specific methodology for constructing descriptive sentences from attributes is also not fully explored. The inherent characteristics or pre-training nuances of the selected LLMs could also influence the quality of extracted features and, consequently, the prediction performance.